# symbol/.style= draw=none, every to/.append style= edge node=node [sloped, allow upside down, auto=false]#$1$

### v1

@@ Abstract.
Inspired by recent work of Aslanyan and Daw, we introduce the notion of $\Sigma$-orbits in the general framework of distinguished categories. In the setting of connected Shimura varieties, this concept contains many instances of (generalized) Hecke orbits from the literature. In the setting of semiabelian varieties, a $\Sigma$-orbit is a subgroup of finite rank. We show that our $\Sigma$-orbits have useful functorial properties and we use them to formulate two general statements of Mordell-Lang type (a weak one and an apparently stronger one). We prove an analogue of a recent theorem of Aslanyan and Daw in this general setting, which we apply to deduce an unconditional result about unlikely intersections in a fibered power of the Legendre family.
@@ 1. Introduction
In this paper, which is inspired by [1], we formulate analogues of the Mordell-Lang conjecture, which we call the strong and the weak Mordell-Lang statement respectively, in a general distinguished category. For this, we introduce the notion of $\Sigma$-orbits. In the setting of semiabelian varieties, $\Sigma$-orbits are subgroups of finite rank and both the strong and the weak Mordell-Lang statement follow from the classical Mordell-Lang conjecture, which is a theorem due to McQuillan [17]. In the setting of connected (mixed) Shimura varieties, we will see that our notion of a $\Sigma$-orbit contains various notions of (generalized) Hecke orbits from the literature. Our weak Mordell-Lang statement for connected pure Shimura varieties implies the Mordell-Lang conjecture for Shimura varieties formulated in [1].
The main result of this article, which is Theorem 7.8, is an analogue of Theorem 4.3 in [1] (see also Remark 7.7) and implies Theorem 1.8 of [2] and an unconditional result (Corollary 8.3) for subvarieties of fibered powers of the Legendre elliptic scheme.
@@
In Section 3, we introduce the notion of $\Sigma$-orbits and prove several basic facts about them. In Section 4, we first describe $\Sigma$-orbits in the semiabelian case and see that they are finite rank subgroups. We then consider the setting of connected pure Shimura varieties and relate the concept of $\Sigma$-orbits with some definitions of (generalized) Hecke orbits appearing in the literature.
Section 5 is devoted to $\Sigma$-special subvarieties, which are weakly special subvarieties coming from a point of a $\Sigma$-orbit: after their definition, we prove a very useful characterization (Lemma 5.4) as well as several important facts. We moreover define $\Sigma$-point-special subvarieties (weakly special subvarieties containing a point of the $\Sigma$-orbit) and show that they are $\Sigma$-special in most examples. In Section 6, we consider the converse problem: we see that $\Sigma$-special subvarieties are $\Sigma$-point-special for almost-split semiabelian varieties and connected pure Shimura varieties, but not in the general setting of semiabelian varieties.
Section 7 starts with the statements of the two above mentioned formulations of the Mordell-Lang statement, a weaker one and an apparently stronger one. We then see that they follow from the Zilber-Pink statement (as defined in Definition 10.1 of [3]) and state and prove our main result, Theorem 7.8. This implies that, in most cases, our weak Mordell-Lang statement implies the strong one.
Finally, in Section 8, we see that the usual Mordell-Lang conjecture for semiabelian varieties (McQuillan’s Theorem [17]) is equivalent to our formulations, we recover Theorem 1.8 of [2], and we deduce an unconditional result (Corollary 8.3) for subvarieties of fibered powers of the Legendre elliptic scheme from our Theorem 7.8 and results of Gao as well as of the second author [13, 7, 8].
For results and conjectures related to our weak Mordell-Lang statement for connected pure Shimura varieties, see [21, 22, 28, 2, 1, 26, 29, 30].
In this article, all fields are of characteristic 00. Varieties and curves are irreducible, subvarieties are closed. We denote by $K$ an algebraically closed field of characteristic 00. Varieties will always be varieties over $K$ if their ground field is not mentioned explicitly.
@@ 2. Preliminaries
    Fiber Dimension - If $\phi:X\to Y$ is a morphism in $\mathfrak{M}$, then $\mathcal{F}(\phi)(\mathcal{F}(X))$ is closed in $\mathcal{F}(Y)$ and there exist morphisms $\phi_{1}:W\to X$, $\phi_{2}:W\to Z$, and $\phi_{3}:Z\to Y$ in $\mathfrak{M}$ such that $\phi\circ\phi_{1}=\phi_{3}\circ\phi_{2}$, $\mathcal{F}(\phi_{1})$ is finite and surjective, $\mathcal{F}(\phi_{3})$ has finite fibers, $\mathcal{F}(\phi_{2})$ is surjective, $\mathcal{F}(\phi_{2})^{-1}(z)$ is irreducible for all $z\in\mathcal{F}(\phi_{2})(\mathcal{F}(W))$, and $\dim_{w}\mathcal{F}(\phi_{2})^{-1}(\mathcal{F}(\phi_{2})(w))$ is a constant function for $w\in\mathcal{F}(W)$.
Note that (A1) and (A4) are slightly stronger than the corresponding axioms in [3] in that we assume that $\chi$ is unique and that $\mathcal{F}(\phi_{2})$ is surjective respectively. All the examples of distinguished categories in [3] also satisfy these stronger versions of (A1) and (A4).
@@
The first part of the lemma is a consequence of (A4). Indeed, in (A4), since $\phi_{1}$ and $\phi_{2}$ are surjective, we have $\phi(X)=\phi_{3}(Z)$. The second part of the lemma follows from the fact that a variety whose underlying topological space is finite must be a point. ∎
@@
In [3], several examples of distinguished categories are given. Most of them also satisfy (A5). In Sections 3, 5, and 7, we always work in a fixed ambient distinguished category unless stated otherwise.
@@ 3. $\Sigma$-orbits
For $X,Z\in\mathfrak{V}$ and $x\in X(K)$ we define the $\Sigma$-orbit $\Sigma_{Z}(x)$ of $x$ in $Z$ to be the set of all $\phi(y)$ where $Y\in\mathfrak{V}$, $\phi:Y\to Z$ is a distinguished morphism, and $y\in Y(K)$ such that there exists a distinguished morphism $\psi:Y\to X$ with finite fibers with $\psi(y)=x$. We call the collection of all $\Sigma_{Z}(x)$, for varying $Z\in\mathfrak{V}$, the $\Sigma$-orbit $\Sigma(x)$ of $x$.
@@
The inclusion $\supset$ is obvious. Let $z^{\prime}\in\chi^{-1}(\Sigma_{Z}(x))$ and set $z=\chi(z^{\prime})$. By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. We consider the fibered product $(X\times Z^{\prime})\times_{X\times Z}Y$ given by the distinguished morphisms $\operatorname{id}_{X}\times\chi:X\times Z^{\prime}\to X\times Z$ and $(\psi,\phi):Y\to X\times Z$. We apply (A2) and Lemma 2.4 to find distinguished morphisms $\phi_{i}:X_{i}\to X\times Z^{\prime}\times Y$, $i=1,\dots,n$, with finite fibers such that $\bigcup_{i=1}^{n}{\phi_{i}(X_{i})}=((X\times Z^{\prime})\times_{X\times Z}Y)_{\operatorname{red}}$. We may assume that $(x,z^{\prime},y)\in\phi_{1}(X_{1}(K))$ and pick $x_{1}\in X_{1}(K)$ such that $\phi_{1}(x_{1})=(x,z^{\prime},y)$. After composing with the appropriate projection, we get distinguished morphisms $\phi^{\prime}:X_{1}\to Z^{\prime}$ and $\psi^{\prime}:X_{1}\to X$, where $\psi^{\prime}$ has finite fibers, such that $\phi^{\prime}(x_{1})=z^{\prime}$ and $\psi^{\prime}(x_{1})=x$. This implies that $z^{\prime}\in\Sigma_{Z^{\prime}}(x)$. ∎
@@
The inclusion $\supset$ is obvious, so suppose that $z\in\Sigma_{Z}(x)$. By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. We consider the fibered product $X^{\prime}\times_{X}Y$. We apply (A2) and Lemma 2.4 to find distinguished morphisms $\phi_{i}:X_{i}\to X^{\prime}\times Y$, $i=1,\dots,n$, with finite fibers such that $\bigcup_{i=1}^{n}{\phi_{i}(X_{i})}=(X^{\prime}\times_{X}Y)_{\operatorname{red}}$. We may assume that $(x^{\prime},y)\in\phi_{1}(X_{1}(K))$ and pick $x_{1}\in X_{1}(K)$ such that $\phi_{1}(x_{1})=(x^{\prime},y)$. After composing with the appropriate projection, we get distinguished morphisms $\psi^{\prime}:X_{1}\to X^{\prime}$ and $\phi^{\prime}:X_{1}\to Y$, both having finite fibers, such that $\psi^{\prime}(x_{1})=x^{\prime}$ and $\phi^{\prime}(x_{1})=y$. It follows that $(\phi\circ\phi^{\prime})(x_{1})=z$. We deduce that $z\in\Sigma_{Z}(x^{\prime})$. ∎
@@
By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. By (A1) the morphism $\chi=(\phi,\psi):Y\to Z\times X$ is distinguished and therefore $\chi(Y)$ is special. Since $\psi$ has finite fibers $Z\times\{x\}\cap\chi(Y)$ is a finite set and thus $\{(z,x)=\chi(y)\}$ is an irreducible component of it. Suppose that (2) holds. By Lemma 2.3, there exists a distinguished morphism $\chi:Y\to Z\times X$ with finite fibers such that $\{(z,x)\}$ is an irreducible component of $\chi(Y)\cap(Z\times\{x\})$. It follows from [3, Lemma 7.4] that the restriction of the projection $Z\times X\to X$ to $\chi(Y)$ has finite fibers. Since $\chi$ has finite fibers, also the induced morphism $\psi:Y\to X$ has finite fibers. It follows that $(z,x)\in\chi(\psi^{-1}(x))\subset\Sigma_{Z\times X}(x)$ and hence $z\in\Sigma_{Z}(x)$. ∎
@@ 4.1. Semiabelian varieties
Let $g\in\Sigma_{G}(x)$. There are distinguished morphisms $\psi:Y\to G$ and $\phi:Y\to G_{0}$ such that $\phi$ has finite fibers and there is a $y\in Y(K)$ with $\psi(y)=g$ and $\phi(y)=x$. First, note that $\phi$ must be surjective since $\phi(Y)$ is a special subvariety of $G_{0}$ that contains $x$. By Proposition 3.3, for any translation $x^{\prime}$ of $x$ by a torsion point of $G_{0}$, we have $\Sigma(x)=\Sigma(x^{\prime})$. Since we also have that $\langle\psi_{1}(x),\dots,\psi_{n}(x)\rangle^{\mathrm{div}}=\langle\psi_{1}(x^{\prime}),\dots,\psi_{n}(x^{\prime})\rangle^{\mathrm{div}}$ for any such $x^{\prime}$, we may and will replace $x$ by some such $x^{\prime}$ and assume without loss of generality that $\phi$ is a homomorphism of algebraic groups. As a surjective homomorphism with finite fibers, $\phi$ is an isogeny. By Lemma 7.3/5 on p. 180 of [4], we may replace $Y$ by $G_{0}$ and $\phi$ by the multiplication by some non-zero integer. Then, $g=t+\sum_{i=1}^{n}a_{i}\psi_{i}(y)$ for some torsion point $t$ and integers $a_{1},\dots,a_{n}$ and $Ny=x$ for some non-zero integer $N$. This clearly implies that $g\in\langle\psi_{1}(x),\dots,\psi_{n}(x)\rangle^{\mathrm{div}}$.
@@
4.2. Connected Shimura varieties
At first glance, our definition of $\Sigma$-orbits in the case of connected pure/mixed Shimura varieties resembles the definition of a generalized Hecke orbit which can be found in Proposition-Definition 3.3 in [24]. In this subsection, we compare these two notions and some other notions of generalized Hecke orbits that appear in some articles.
Slightly different definitions of (connected) pure/mixed Shimura data are used in these papers. We will always use the definition from [3] and translate all definitions of generalized Hecke orbits to this setting.
In the following, we freely use terminology from Section 3 of [3]. Let $(P,X^{+},\Gamma)$ be an object of the distinguished category $\mathfrak{C}_{\mathrm{pSv}}$ of connected (pure) Shimura varieties with associated functor $\mathcal{F}_{\mathrm{pSv}}$. Set $S=\mathcal{F}_{\mathrm{pSv}}(P,X^{+},\Gamma)$ and let $x\in S(\mathbb{C})$. There is a canonical identification of $S(\mathbb{C})$ with $\Gamma\backslash X^{+}$. Let $\widetilde{x}$ be a lift of $x$ to $X^{+}\subset\mathrm{Hom}(\mathbb{S}_{\mathbb{C}},P_{\mathbb{C}})$, where $\mathbb{S}$ denotes the Deligne torus, let $M$ denote the smallest algebraic subgroup of $P$ (defined over $\mathbb{Q}$) such that the image of $\widetilde{x}$ is contained in $M_{\mathbb{C}}$, and let $\widetilde{x}_{0}:\mathbb{S}_{\mathbb{C}}\to M_{\mathbb{C}}$ denote the induced homomorphism. Since $M$ is automatically connected, it follows from Lemma 4.3 in [3] that there is a Shimura subdatum $(M,X_{M}^{+})$ of $(P,X^{+})$ such that $\widetilde{x}\in X_{M}^{+}$. Thanks to Lemma 2.2 in [21] and Proposition 2.2 in [27], we find an object $(M,X_{M}^{+},\Gamma_{M})$ of $\mathfrak{C}_{\mathrm{pSv}}$ and a morphism $(M,X_{M}^{+},\Gamma_{M})\to(P,X^{+},\Gamma)$ of $\mathfrak{C}_{\mathrm{pSv}}$ extending the Shimura morphism $(M,X_{M}^{+})\to(P,X^{+})$.

### v2

# Hecke orbits and the Mordell-Lang conjecture in distinguished categories

@@ Abstract.
Inspired by recent work of Aslanyan and Daw, we introduce the notion of $\Sigma$-orbits in the general framework of distinguished categories. In the setting of connected Shimura varieties, this concept contains many instances of (generalized) Hecke orbits from the literature. In the setting of semiabelian varieties, a $\Sigma$-orbit is a subgroup of finite rank. We show that our $\Sigma$-orbits have useful functorial properties and we use them to formulate two general statements of Mordell-Lang type (one of them implying the other one). We prove an analogue of a recent theorem of Aslanyan and Daw in this general setting, which we apply to deduce an unconditional result about unlikely intersections in a fibered power of the Legendre family. In an appendix, we prove an unconditional Zilber-Pink result for subvarieties of $\mathcal{A}_{g}$ that cannot be defined over the algebraic numbers.
@@ 1. Introduction
In this paper, which is inspired by [1], we formulate analogues of the Mordell-Lang conjecture, which we call the strong and the weak Mordell-Lang statement respectively, in a general distinguished category. For this, we introduce the notion of $\Sigma$-orbits. In the setting of semiabelian varieties, $\Sigma$-orbits are subgroups of finite rank and both the strong and the weak Mordell-Lang statement follow from the classical Mordell-Lang conjecture, which is a theorem due to McQuillan [20]. In the setting of connected pure Shimura varieties, we will see that our notion of a $\Sigma$-orbit contains various notions of (generalized) Hecke orbits from the literature. In particular, we will see in Subsection 4.2 that, for connected pure Shimura varieties, our notion of a $\Sigma$-orbit is equivalent to the notion of hybrid Hecke orbit introduced in [31]. Our weak Mordell-Lang statement for connected pure Shimura varieties implies the Mordell-Lang conjecture for Shimura varieties formulated in [1]; for connected pure Shimura varieties of abelian type, our weak Mordell-Lang statement has in the meantime been proved in [33].
The main result of this article, which is Theorem 7.9, is an analogue of Theorem 4.3 in [1] (see also Remark 7.8). Combined with results from the literature, Theorem 7.9 implies Theorem 1.8 of [2] and unconditional results for subvarieties of connected pure Shimura varieties of abelian type (Corollary 8.2) as well as for subvarieties of fibered powers of the Legendre elliptic scheme (Corollary 8.5).
Thanks to a discussion at the third JNT biennial conference, we realized that the methods employed in [4] and the present article also lead to a resolution of some cases of the Zilber-Pink conjecture for connected pure Shimura varieties with simple adjoint group. We include the proof of the relevant statement about distinguished categories in Appendix A. The following theorem is a special case of Theorem A.2 (see Remark A.3). We use $\mathcal{A}_{g}$ to denote the coarse moduli space of principally polarized abelian varieties of dimension $g$ over $\bar{\mathbb{Q}}$ ($g\in\mathbb{N}$).
Theorem 1.1.
Let $g\in\mathbb{N}$ and let $K$ be an algebraically closed field of characteristic 00. If $V\subset(\mathcal{A}_{g})_{K}$ is an irreducible closed subvariety that is not the base change of a closed subvariety of $\mathcal{A}_{g}$ (in particular, $K\neq\bar{\mathbb{Q}}$) and is not contained in a proper special subvariety of $(\mathcal{A}_{g})_{K}$, we have that the union of the intersections of $V$ with all special subvarieties of $(\mathcal{A}_{g})_{K}$ of codimension $>\dim V$ is not Zariski dense in $V$.
If $g=1$, Theorem 1.1 holds trivially while, if $g=2$ and $\dim V=1$, we have proved it in [4] (if $g=2$ and $\dim V=2$, Theorem 1.1 again holds trivially since special points are algebraic).
@@
After recalling some preliminaries in Section 2, we introduce in Section 3 the notion of $\Sigma$-orbits and prove several basic facts about them. In Section 4, we first describe $\Sigma$-orbits in the semiabelian case and see that they are finite rank subgroups. We then consider the setting of connected pure Shimura varieties and relate the concept of $\Sigma$-orbits with some definitions of (generalized) Hecke orbits appearing in the literature.
Section 5 is devoted to $\Sigma$-special subvarieties, which are weakly special subvarieties coming from a point of a $\Sigma$-orbit $\Sigma$: after their definition, we prove a very useful characterization (Lemma 5.4) as well as several important facts. We moreover define $\Sigma_{Z}$-point-special subvarieties of a distinguished variety $Z$ (weakly special subvarieties containing a point of $\Sigma$) and show that they are $\Sigma$-special in most examples. In Section 6, we consider the converse problem: we show that $\Sigma$-special subvarieties are $\Sigma_{Z}$-point-special for almost-split semiabelian varieties and connected pure Shimura varieties, but not in the general setting of semiabelian varieties.
Section 7 starts with the statements of the two above mentioned formulations of the Mordell-Lang statement, one of them implying the other one. We then show that they follow from the Zilber-Pink statement (as defined in Definition 10.1 of [4]) and state and prove our main result, Theorem 7.9.
Finally, in Section 8, we show that the usual Mordell-Lang conjecture for semiabelian varieties (McQuillan’s Theorem [20]) is equivalent to our formulations, we recover Theorem 1.8 of [2], and we deduce an unconditional result (Corollary 8.2) for connected pure Shimura varieties of abelian type from our Theorem 7.9 together with a recent result of Richard and Yafaev [33]. We also deduce an unconditional result (Corollary 8.5) for subvarieties of fibered powers of the Legendre elliptic scheme from our Theorem 7.9 and results of Gao as well as of the second-named author [16, 9, 10].
For results and conjectures related to our weak Mordell-Lang statement for connected pure Shimura varieties, see [24, 25, 30, 2, 28, 34, 32, 1, 31, 33].
In this article, all fields are of characteristic 00. Varieties and curves are irreducible, subvarieties are closed. While (mixed) Shimura varieties as usually defined might not be irreducible, we consider here only connected (mixed) Shimura varieties, which are irreducible. We denote by $K$ an algebraically closed field of characteristic 00. Varieties will always be varieties over $K$ if their ground field is not mentioned explicitly.
@@ 2. Preliminaries
    Fiber Dimension - If $\phi:X\to Y$ is a morphism in $\mathfrak{M}$, then $\mathcal{F}(\phi)(\mathcal{F}(X))$ is closed in $\mathcal{F}(Y)$ and there exist morphisms $\phi_{1}:W\to X$, $\phi_{2}:W\to Z$, and $\phi_{3}:Z\to Y$ in $\mathfrak{M}$ such that $\phi\circ\phi_{1}=\phi_{3}\circ\phi_{2}$, $\mathcal{F}(\phi_{1})$ is finite and surjective, $\mathcal{F}(\phi_{3})$ has finite fibers, $\mathcal{F}(\phi_{2})$ is surjective, $\mathcal{F}(\phi_{2})^{-1}(z)$ is irreducible for all $z\in\mathcal{F}(Z)$, and $\dim_{w}\mathcal{F}(\phi_{2})^{-1}(\mathcal{F}(\phi_{2})(w))$ is a constant function for $w\in\mathcal{F}(W)$.
@@
Note that (A4) implies that $\dim_{x}\phi^{-1}(\phi(x))$ is a constant function for $x\in X$ (see Section 2 of [4]). If $L$ is an algebraically closed field that contains $K$ and $\mathfrak{C}$ is a distinguished category over $K$, then we obtain a distinguished category $\mathfrak{C}_{L}$ over $L$ by composing $\mathcal{F}$ with the base change functor from $K$ to $L$.
@@
The first part of the lemma is a consequence of (A4). Indeed, in (A4), since $\phi_{1}$ and $\phi_{2}$ are surjective, we have $\phi(X)=\phi_{3}(Z)$.
The second part of the lemma follows from the fact that a variety whose underlying topological space is finite must be a point. ∎
@@
In [4], several examples of distinguished categories are given. Most of them also satisfy (A5). In particular, semiabelian varieties, connected pure Shimura varieties, and connected mixed Shimura varieties of Kuga type all form distinguished categories that satisfy (A5) and to which we will apply the results of this article. In Sections 3, 5, and 7, we always work in a fixed ambient distinguished category unless stated otherwise.
@@ 3. $\Sigma$-orbits
For $X,Z\in\mathfrak{V}$ and $x\in X(K)$ we define the $\Sigma$-orbit $\Sigma_{Z}(x)$ of $x$ in $Z$ to be the set of all $\phi(y)$ where $Y\in\mathfrak{V}$, $\phi:Y\to Z$ is a distinguished morphism, and $y\in Y(K)$ such that there exists a distinguished morphism $\psi:Y\to X$ with finite fibers with $\psi(y)=x$.
We call the collection of all $\Sigma_{Z}(x)$, for varying $Z\in\mathfrak{V}$, the $\Sigma$-orbit $\Sigma(x)$ of $x$.
@@
The inclusion $\supset$ is obvious. Let $z^{\prime}\in\chi^{-1}(\Sigma_{Z}(x))$ and set $z=\chi(z^{\prime})$. By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. We consider the fibered product $(X\times Z^{\prime})\times_{X\times Z}Y$ given by the distinguished morphisms $\operatorname{id}_{X}\times\chi:X\times Z^{\prime}\to X\times Z$ and $(\psi,\phi):Y\to X\times Z$.
We apply (A2) and Lemma 2.4 to find distinguished morphisms $\phi_{i}:X_{i}\to X\times Z^{\prime}\times Y$, $i=1,\dots,n$, with finite fibers such that $\bigcup_{i=1}^{n}{\phi_{i}(X_{i})}=((X\times Z^{\prime})\times_{X\times Z}Y)_{\operatorname{red}}$. We may assume that $(x,z^{\prime},y)\in\phi_{1}(X_{1}(K))$ and pick $x_{1}\in X_{1}(K)$ such that $\phi_{1}(x_{1})=(x,z^{\prime},y)$. After composing with the appropriate projection, we get distinguished morphisms $\phi^{\prime}:X_{1}\to Z^{\prime}$ and $\psi^{\prime}:X_{1}\to X$, where $\psi^{\prime}$ has finite fibers, such that $\phi^{\prime}(x_{1})=z^{\prime}$ and $\psi^{\prime}(x_{1})=x$. This implies that $z^{\prime}\in\Sigma_{Z^{\prime}}(x)$. ∎
@@
The inclusion $\supset$ is obvious, so suppose that $z\in\Sigma_{Z}(x)$. By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. We consider the fibered product $X^{\prime}\times_{X}Y$.
We apply (A2) and Lemma 2.4 to find distinguished morphisms $\phi_{i}:X_{i}\to X^{\prime}\times Y$, $i=1,\dots,n$, with finite fibers such that $\bigcup_{i=1}^{n}{\phi_{i}(X_{i})}=(X^{\prime}\times_{X}Y)_{\operatorname{red}}$. We may assume that $(x^{\prime},y)\in\phi_{1}(X_{1}(K))$ and pick $x_{1}\in X_{1}(K)$ such that $\phi_{1}(x_{1})=(x^{\prime},y)$. After composing with the appropriate projection, we get distinguished morphisms $\psi^{\prime}:X_{1}\to X^{\prime}$ and $\phi^{\prime}:X_{1}\to Y$, both having finite fibers, such that $\psi^{\prime}(x_{1})=x^{\prime}$ and $\phi^{\prime}(x_{1})=y$. It follows that $(\phi\circ\phi^{\prime})(x_{1})=z$. We deduce that $z\in\Sigma_{Z}(x^{\prime})$. ∎
@@
By definition there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$, where the latter has finite fibers, and there is $y\in Y(K)$ such that $\psi(y)=x$ and $\phi(y)=z$. By (A1) the morphism $\chi=(\phi,\psi):Y\to Z\times X$ is distinguished and therefore $\chi(Y)$ is special. Since $\psi$ has finite fibers $Z\times\{x\}\cap\chi(Y)$ is a finite set and thus $\{(z,x)=\chi(y)\}$ is an irreducible component of it.
Suppose that (2) holds. By Lemma 2.3, there exists a distinguished morphism $\chi:Y\to Z\times X$ with finite fibers such that $\{(z,x)\}$ is an irreducible component of $\chi(Y)\cap(Z\times\{x\})$.
It follows from [4, Lemma 7.4] that the restriction of the projection $Z\times X\to X$ to $\chi(Y)$ has finite fibers. Since $\chi$ has finite fibers, also the induced morphism $\psi:Y\to X$ has finite fibers. It follows that $(z,x)\in\chi(\psi^{-1}(x))\subset\Sigma_{Z\times X}(x)$ and hence $z\in\Sigma_{Z}(x)$. ∎
@@ 4.1. Semiabelian varieties
We first show $\subset$”. Let $g\in\Sigma_{G}(x)$. There are distinguished morphisms $\psi:Y\to G$ and $\phi:Y\to G_{0}$ such that $\phi$ has finite fibers and there is a $y\in Y(K)$ with $\psi(y)=g$ and $\phi(y)=x$. First, note that $\phi$ must be surjective since $\phi(Y)$ is a special subvariety of $G_{0}$ that contains $x$. By Proposition 3.3, for any translate $x^{\prime}$ of $x$ by a torsion point of $G_{0}$, we have $\Sigma(x)=\Sigma(x^{\prime})$. Since we also have that $\langle\psi_{1}(x),\dots,\psi_{n}(x)\rangle^{\mathrm{div}}=\langle\psi_{1}(x^{\prime}),\dots,\psi_{n}(x^{\prime})\rangle^{\mathrm{div}}$ for any such $x^{\prime}$, we may and will replace $x$ by some such $x^{\prime}$ and assume without loss of generality that $\phi$ is a homomorphism of algebraic groups. As a surjective homomorphism with finite fibers, $\phi$ is an isogeny. By Lemma 7.3/5 on p. 180 of [6], we may replace $Y$ by $G_{0}$ and $\phi$ by the multiplication by some non-zero integer. Then, $g=t+\sum_{i=1}^{n}a_{i}\psi_{i}(y)$ for some torsion point $t$ and integers $a_{1},\dots,a_{n}$ and $Ny=x$ for some non-zero integer $N$. This clearly implies that $g\in\langle\psi_{1}(x),\dots,\psi_{n}(x)\rangle^{\mathrm{div}}$.
@@
4.2. Connected pure Shimura varieties
At first glance, our definition of $\Sigma$-orbits in the case of connected pure/mixed Shimura varieties resembles the definition of a generalized Hecke orbit which can be found in Proposition-Definition 3.3 in [27]. In this subsection, we compare these two notions and some other notions of generalized Hecke orbits that appear throughout the literature.
Several slightly different definitions of (connected) pure/mixed Shimura data are used in the literature. We will always use the definition from [4] and translate all definitions of generalized Hecke orbits to this setting.
In the following, we freely use terminology from Section 3 of [4]. Let $(P,X^{+},\Gamma)$ be an object of the distinguished category $\mathfrak{C}_{\mathrm{pSv}}$ of connected pure Shimura varieties with associated functor $\mathcal{F}_{\mathrm{pSv}}$. Set $S=\mathcal{F}_{\mathrm{pSv}}(P,X^{+},\Gamma)$ and let $x\in S(\mathbb{C})$. There is a canonical identification of $S(\mathbb{C})$ with $\Gamma\backslash X^{+}$. Let $\widetilde{x}$ be a lift of $x$ to $X^{+}\subset\mathrm{Hom}(\mathbb{S}_{\mathbb{C}},P_{\mathbb{C}})$, where $\mathbb{S}$ denotes the Deligne torus, let $M$ denote the smallest algebraic subgroup of $P$ (defined over $\mathbb{Q}$) such that the image of $\widetilde{x}$ is contained in $M_{\mathbb{C}}$, and let $\widetilde{x}_{0}:\mathbb{S}_{\mathbb{C}}\to M_{\mathbb{C}}$ denote the induced homomorphism. Since $M$ is automatically connected, it follows from Lemma 4.3 in [4] that there is a Shimura subdatum $(M,X_{M}^{+})$ of $(P,X^{+})$ such that $\widetilde{x}_{0}\in X_{M}^{+}$. Thanks to Lemma 2.2 in [24] and Proposition 2.2 in [29], we find an object $(M,X_{M}^{+},\Gamma_{M})$ of $\mathfrak{C}_{\mathrm{pSv}}$ and a morphism $(M,X_{M}^{+},\Gamma_{M})\to(P,X^{+},\Gamma)$ of $\mathfrak{C}_{\mathrm{pSv}}$ extending the Shimura morphism $(M,X_{M}^{+})\to(P,X^{+})$.